OpenStudy (anonymous):

Find an inverse for 43 modulo 660. That is, find an integer s such that 43s=1(mod 660)

jimthompson5910 (jim_thompson5910):

Are you familiar with the extended euclidean algorithm?

OpenStudy (anonymous):

umm kinda

jimthompson5910 (jim_thompson5910):

First start with the idea that 43s = 1 (mod 660) This is basically saying that 660 divides (or is a factor of 43s-1 So 660 | (43s - 1)

jimthompson5910 (jim_thompson5910):

When we say 660 | (43s - 1), we mean that there is an integer k such that 660k = 43s - 1

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

Now move the 43s term over and make that -1 positive 660k = 43s - 1 660k - 43s = -1 43s - 660k = 1

jimthompson5910 (jim_thompson5910):

We now apply the euclidean algorithm: 660 = 15*43 + 15 43 = 2*15 + 13 15 = 1*13 + 2 13 = 6*2 + 1 This shows us that the GCD between the two is 1, but more importantly, is we can rewrite that last term of 1 in terms of 660 and 43

jimthompson5910 (jim_thompson5910):

Now take each equation and isolate the remainders: 660 = 15*43 + 15 ------> 15 = 660 - 15*43 43 = 2*15 + 13 ------> 13 = 43 - 2*15 15 = 1*13 + 2 ------> 2 = 15 - 1*13 13 = 6*2 + 1 ------> 1 = 13 - 6*2

jimthompson5910 (jim_thompson5910):

With me so far?

OpenStudy (anonymous):

yesssss

jimthompson5910 (jim_thompson5910):

Take 1 = 13 - 6*2 and replace the "2" with 15 - 1*13 We can do this because 2 = 15 - 1*13 So 1 = 13 - 6*2 1 = 13 - 6*(15 - 1*13) 1 = 13 - 6*15 + 6*13

jimthompson5910 (jim_thompson5910):

Then simplify to get 1 = -6*15 + 7*13

OpenStudy (anonymous):

okkk got that

jimthompson5910 (jim_thompson5910):

Then replace each copy of "13" with 43 - 2*15 (this is possible since 13 = 43 - 2*15) 1 = -6*15 + 7*(43 - 2*15) 1 = -6*15 + 7*43 - 14*15 1 = 7*43 - 20*15

OpenStudy (anonymous):

okkkkkkkkkkkkkk

jimthompson5910 (jim_thompson5910):

Finally, replace "15" with 660 - 15*43 to get 1 = 7*43 - 20*15 1 = 7*43 - 20*(660 - 15*43) 1 = 7*43 - 20*660 + 300*43 1 = 307*43 - 20*660 So 307*43 - 20*660 = 1

OpenStudy (anonymous):

okkkkkkkk

jimthompson5910 (jim_thompson5910):

Now get it back into the original form we had it 307*43 - 20*660 = 1 -20*660 = 1 - 307*43 20*660 = 307*43 - 1 20*660 = 43*307 - 1 660 | (43*307 - 1) 43*307 = 1 (mod 660) Therefore, s = 307 which tells us that the multiplicative inverse of 43 (mod 660) is 307 (mod 660)

OpenStudy (anonymous):

okkkk yay

OpenStudy (anonymous):

THANKSSSSSSSSS

jimthompson5910 (jim_thompson5910):

you're welcome